1. Explain the difference between testing a single mean and testing the difference between two means.


2. What two assumptions must be met when you are using the z test to test differences between two means? Can the sample standard deviations 𝑠1 and 𝑠2 be used in place of the population standard deviations 𝜎1 and 𝜎2?Β 


1
Expert's answer
2021-10-18T12:57:14-0400

1.

for z- test:

testing a single mean:

"H_0:\mu=A"

"H_a:\mu\neq A"

"z=\frac{x-\mu}{\sigma}"


testing the difference between two means:

"H_0:\mu_1=\mu_2"

"H_a:\mu_1\neq \mu_2"

"z=\frac{\mu_1-\mu_2}{\sqrt{\sigma_1^2/n_1+\sigma^2_2/n_2}}"


2.

It can be:

"\sigma_1=\sigma_2" or "\sigma_1\neq\sigma_2"


The sample standard deviations 𝑠1 and 𝑠2 cannot be used in place of the population standard deviations 𝜎1 and 𝜎2.

"s_1=\sigma_1/\sqrt{n_1}, s_2=\sigma_2/\sqrt{n_2}"